Question 13 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Which country has a comparative advantage growing grapes? Spain France Shi Grapes (in tons) 75 Shi Grapes (in tons 50 40 10 20 15 45 75 45 30 40 50 20 10
Answer #1

Comparative Advantage

A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

Spain: It can either produce 75 tons of grapes or 50 ships.

75 tons of grapes = 50 Ships

1 ton of grapes = 50/75 = 0.67 ships

Opportunity cost of 1 ton of grapes is 0.67 ships

France:

50 tons of grapes = 75 Ships

1 ton of grapes = 75/50 = 1.5 ships

Opportunity cost of 1 ton of grapes is 1.5 ships

Therefore, Spain has comparative advantage in the production of grapes.